# Geometric Hardy inequalities on the Heisenberg groups via convexity

### v2

@@ 3.2. Hardy inequalities with respect to the gauge pseudodistance
$\int_{A_{k}}\nabla_{H}d_{k}\cdot\nabla_{H}\phi_{\epsilon}\,{\rm d}\xi=$ $-\int_{A_{k}}\phi_{\epsilon}\,\Delta_{H}d_{k}\,{\rm d}\xi+\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{k}\cdot\nu_{k,H}\,dS$
$\geq$ $\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{k}\cdot\nu_{k,H}\,dS.$
@@
(31) $\int_{\Omega}\nabla_{H}d\cdot\nabla_{H}\phi_{\epsilon}\,{\rm d}\xi\geq\sum_{k=1}^{m}\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{k}\cdot\nu_{k,H}\,dS.$
Now, each boundary $\partial A_{k}$ consists of outer parts where $\phi$ vanishes as well as of common boundaries with other sets $A_{j}$, $j\neq k$. Let us fix $k,j$ such a set $A_{k}$ and $A_{j}$ share such a common boundary $S_{kj}$. Letting, as above, $\nu_{H}=(\nu_{x}+2\nu_{t}y,\nu_{y}-2\nu_{t}x)$ where $\nu$ is the normal vector on $S_{kj}$ which is outer with respect to $A_{k}$, we conclude that the two contributions on the surface $S_{kj}$ from $A_{k}$ and $A_{j}$ add up to
[…]
The surface $S_{kj}$ is a level set for the function $d_{k}-d_{j}$ and at each point $\xi\in S_{kj}$ there holds $\nabla d_{k}-\nabla d_{j}=\lambda\nu$ where $\lambda=\lambda(\xi)\geq 0$. We therefore have on $S_{kj}$
$(\nabla_{H}\,d_{k}-\nabla_{H}\,d_{j})\cdot\nu_{H}$
$=\big{(}\nabla_{\!x}\,d_{k}-\nabla_{\!x}\,d_{j}+2d_{k,t}y-2d_{j,t}y,\nabla_{\!y}\,d_{k}-\nabla_{\!y}\,d_{j}-2d_{k,t}x+2d_{j,t}x\big{)}$
$\quad\quad\cdot\big{(}\nu_{x}+2y\nu_{t},\nu_{y}-2x\nu_{t}\big{)}$
@@ 3.3. Hardy inequalities with respect to the Carnot-Carathéodory distance
The proof follows using standard arguments by Lemma 13, since by (42) $t\rightarrow 0$ implies that $\phi\rightarrow 0$. ∎
@@
For simplicity we write $d$ instead of $d_{\rho}$. $({\rm i})$ We first note that by invariance under group action we may assume that $\Pi=\Pi_{0}$. By Lemma 13 the distance to the boundary of a point $\xi=(x,y,t)$, $(x,y)\neq(0,0)$, is given by
$d(r,t)=\frac{\phi}{\cos\phi}r=:B(\phi)r.$
We will see below that $d_{r}(r,t)\leq 0$. Therefore, we have
(44) $\Delta_{{\mathbb{H}}^{n}}d=$ $\;d_{rr}+\frac{2n-1}{r}d_{r}+4r^{2}d_{tt}$
$\leq$ $\;d_{rr}+\frac{1}{r}d_{r}+4r^{2}d_{tt}$
@@
(46) $\Delta_{{\mathbb{H}}^{n}}d\leq\frac{1}{r(Q^{\prime})^{3}}\bigg{(}4(Q^{2}+1)(B^{\prime\prime}Q^{\prime}-Q^{\prime\prime}B^{\prime})+B\,(Q^{\prime})^{3}\bigg{)}.$
@@
(ii) The required Hardy inequality will follow by applying Theorem 1 provided we establish that $\Delta_{p,{\mathbb{H}}^{n}}d\leq 0$ in the distributional sense in $\Pi$. Now, by [MSC01, Theorem 3.1] (or by a direct computation for $r\neq 0$) we have $|\nabla_{{\mathbb{H}}^{n}}d|=1$ a.e.
Hence the above condition can be simplified to $\Delta_{{\mathbb{H}}^{n}}d\leq 0$ which is in particular independent of $p>1$; see also [BFT04]. So the result follows from (i). The sharpness of the constant follows from Theorem 1 (b), Proposition 14 and the fact that $|\nabla_{{\mathbb{H}}^{n}}d_{\rho}|=1$ a.e., cf. [MSC01]. ∎
@@
$(1b)$ Let $t>0$. The point $(0,0,t)$ has as nearest boundary points all points $(x^{\prime},y^{\prime},0)$ with $x^{\prime 2}+y^{\prime 2}=2t/\pi$.
@@
(2a) Assume that $\xi=(x,y,t)\in\Pi_{0}$ with $(x,y)\neq(x^{\prime},y^{\prime})$ has $\xi^{\prime}$ as its nearest boundary point and that $d(\xi)=\rho$. By (47) $\phi=\rho/r^{\prime}$ and the point $\xi$ is now uniquely determined by (37) and (42). It is now easy to see that this point $\xi$ has indeed $\xi^{\prime}$ as its nearest boundary point and $d_{\rho}(\xi)=\rho$.
@@
For simplicity we write $d$ instead of $d_{\rho}$. Let $u\in{C^{\infty}_{c}}(\Omega)$ be fixed. We consider a bounded convex polytope $\Omega^{\prime}$ such that
@@
Let us denote by $d^{\prime}$ the Carnot-Carathéodory distance to $\partial\Omega^{\prime}$. Combining Theorem 1 and Lemma 17 we obtain that
$\int_{\Omega^{\prime}}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{\Omega^{\prime}}\frac{|u|^{p}}{d^{\prime p}}{\rm d}\xi\;.$
Hence, since $d^{\prime}\leq d$ in $\Omega^{\prime}$,
$\int_{\Omega}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi=\int_{\Omega^{\prime}}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{\Omega^{\prime}}\frac{|u|^{p}}{d^{\prime p}}{\rm d}\xi\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{\Omega}\frac{|u|^{p}}{d^{p}}{\rm d}\xi$
@@
We point that the fact that the key property that $d^{\prime}\leq d$, for $d$ and $d^{\prime}$ as in the proof of Theorem 18 in the bounded convex polytope $\Omega^{\prime}$ and reflects the geometric nature of the Carnot-Carathéodory distance, meaning in particular that that the latter is a “true” distance respecting the geometry of ${\mathbb{H}}^{n}$.
@@ 4. Hardy inequalities on stratified groups of step two
[…]
@@
where $\langle Bg^{(1)},g^{\prime(1)}\rangle$ denotes the $(n-m)$-tuple
@@ 4.1. On the distance function from the boundary of bounded convex domains in stratified groups
Let $\mathbb{H}^{n}$ be the Heisenberg group and $\Pi_{0}=\{(x,y,t)\in\mathbb{H}^{n}:t>0\}$. We shall prove that the the distance $d_{N}$ (cf. (13)) is not necessarily weakly $H$-concave.
@@ 4.2. Hardy inequalities with respect to the Euclidean distance
In this section we prove that the Euclidean distance to the boundary on a convex, bounded domain $\Omega$ is weakly $H$-concave and superharmonic. This provides an alternative proof of the $L^{2}$-Hardy inequality for such domains.
@@
From Theorem 22 we immediately obtain the following result; the sharpness of the constant 1/4 follows under the hypotheses of part (b) of Theorem 1.
@@
Let $G$ be a stratified group of step two and let $\Omega\subset G$ be a bounded domain which is convex in the Euclidean sense. Then
${\rm(i)}$ $\quad\Delta_{H}d\leq 0\;\mbox{ in the distributional sense in }\Omega;$
@@
$\hskip 42.67912pt\int_{\Omega}|\nabla_{H}u|^{2}{\rm d}g\geq\frac{1}{4}\int_{\Omega}\frac{|\nabla_{H}d|^{2}}{d^{2}}u^{2}{\rm d}g\;,\quad u\in{C^{\infty}_{c}}(\Omega),$
@@
By Theorem 22 the distance function is weakly $H$-concave. Let $X_{1},\ldots,$ $X_{m}$ be the vector fields that generate the first stratum $V_{1}$ of the corresponding Lie algebra. By Theorem [DGN03, Theorem 8.1] each $X_{k}^{2}d$, $k=1,\ldots,m$, is a non-positive Radon measure on $\Omega$; this proves (i). Part (ii) now follows from Theorem 1. ∎

### v3

@@ 3.2. Hardy inequalities with respect to the gauge pseudodistance
$\int_{A_{k}}\nabla_{H}d_{N,k}\cdot\nabla_{H}\phi_{\epsilon}\,{\rm d}\xi=$ $-\int_{A_{k}}\phi_{\epsilon}\,\Delta_{H}d_{N,k}\,{\rm d}\xi+\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{N,k}\cdot\nu_{k,H}\,dS$
$\geq$ $\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{N,k}\cdot\nu_{k,H}\,dS.$
@@
(34) $\int_{\Omega}\nabla_{H}d_{N}\cdot\nabla_{H}\phi_{\epsilon}\,{\rm d}\xi\geq\sum_{k=1}^{m}\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{N,k}\cdot\nu_{k,H}\,dS.$
Now, each boundary $\partial A_{k}$ consists of outer parts where $\phi$ vanishes as well as of common boundaries with other sets $A_{j}$, $j\neq k$. Let us fix $k,j$ such a set $A_{k}$ and $A_{j}$ share such a common boundary $S_{kj}$. Let $\nu$ be the normal vector on $S_{kj}$ which is outer with respect to $A_{k}$ (the existence of $\nu$ is shown in the Appendix in case where the two hyperplanes are not vertical; otherwise one may use an approximation argument, see also [Lar16]). Defining $\nu_{H}=(\nu_{x}+2\nu_{t}y,\nu_{y}-2\nu_{t}x)$ we conclude that the two contributions on the surface $S_{kj}$ from $A_{k}$ and $A_{j}$ add up to
$\int_{S_{kj}}\!\!\phi_{\epsilon}\,(\nabla_{H}d_{N,k}-\nabla_{H}d_{N,j})\cdot\nu_{H}dS.$
The surface $S_{kj}$ is a level set for the function $d_{N,k}-d_{N,j}$ and at each point $\xi\in S_{kj}$ there holds $\nabla d_{N,k}-\nabla d_{N,j}=\lambda\nu$ where $\lambda=\lambda(\xi)\geq 0$. We therefore have on $S_{kj}$
$(\nabla_{H}\,d_{N,k}-\nabla_{H}\,d_{N,j})\cdot\nu_{H}=\big(\nabla_{\!x}\,d_{N,k}-\nabla_{\!x}\,d_{N,j}+2(d_{N,k})_{t}y-2(d_{N,j})_{t}y,$
$\quad\quad\nabla_{\!y}\,d_{N,k}-\nabla_{\!y}\,d_{N,j}-2(d_{N,k})_{t}x+2(d_{N,j})_{t}x\big)\cdot\big(\nu_{x}+2y\nu_{t},\nu_{y}-2x\nu_{t}\big)$
@@ 3.3. Hardy inequalities with respect to the Carnot-Carathéodory distance
The proof follows from Lemma 14 using standard arguments, since by (45) $t\rightarrow 0$ implies that $\phi\rightarrow 0$. ∎
@@
$({\rm i})$ We first note that by invariance under group action we may assume that $\Pi=\Pi_{0}$. By Lemma 14 the distance to the boundary of a point $\xi=(x,y,t)$, $(x,y)\neq(0,0)$, is given by
$d_{\rho}(r,t)=\frac{\phi}{\cos\phi}r=:B(\phi)r.$
We will see below that $d_{\rho,r}(r,t)\leq 0$. Therefore, we have
(47) $\Delta_{{\mathbb{H}}^{n}}d_{\rho}=$ $\;d_{\rho,rr}+\frac{2n-1}{r}d_{\rho,r}+4r^{2}d_{\rho,tt}$
$\leq$ $\;d_{\rho,rr}+\frac{1}{r}d_{\rho,r}+4r^{2}d_{\rho,tt}$
@@
(49) $\Delta_{{\mathbb{H}}^{n}}d_{\rho}\leq\frac{1}{r(Q^{\prime})^{3}}\bigg(4(Q^{2}+1)(B^{\prime\prime}Q^{\prime}-Q^{\prime\prime}B^{\prime})+B\,(Q^{\prime})^{3}\bigg).$
@@
(ii) The required Hardy inequality will follow by applying Theorem 1 provided we establish that $\Delta_{p,{\mathbb{H}}^{n}}d_{\rho}\leq 0$ in the distributional sense in $\Pi$. Now, by [MSC01, Theorem 3.1] (or by a direct computation for $r\neq 0$) we have $|\nabla_{{\mathbb{H}}^{n}}d_{\rho}|=1$ a.e.
Hence the above condition can be simplified to $\Delta_{{\mathbb{H}}^{n}}d_{\rho}\leq 0$ which is in particular independent of $p>1$; see also [BFT04]. So the result follows from (i). The sharpness of the constant follows from Theorem 1 (b), Proposition 15 and the fact that $|\nabla_{{\mathbb{H}}^{n}}d_{\rho}|=1$ a.e., cf. [MSC01]. ∎
@@
$(1b)$ Let $t>0$. The point $(0,0,t)$ has as nearest boundary points all points $(x^{\prime},y^{\prime},0)$ with $|x^{\prime}|^{2}+|y^{\prime}|^{2}=2t/\pi$.
@@
(2a) Assume that $\xi=(x,y,t)\in\Pi_{0}$ with $(x,y)\neq(x^{\prime},y^{\prime})$ has $\xi^{\prime}$ as its nearest boundary point and that $d_{\rho}(\xi)=\rho$. By (50) $\phi=\rho/r^{\prime}$ and the point $\xi$ is now uniquely determined by (40) and (45). It is now easy to see that this point $\xi$ has indeed $\xi^{\prime}$ as its nearest boundary point and $d_{\rho}(\xi)=\rho$.
@@
Let $u\in{C^{\infty}_{c}}(\Omega)$ be fixed. We consider a bounded convex polytope $\Omega^{\prime}$ such that
@@
Let us denote by $d_{\rho}^{\prime}$ the Carnot-Carathéodory distance to $\partial\Omega^{\prime}$. Combining Theorem 1 and Lemma 18 we obtain that
$\int_{\Omega^{\prime}}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big(\frac{p-1}{p}\Big)^{p}\int_{\Omega^{\prime}}\frac{|u|^{p}}{d_{\rho}^{\prime p}}{\rm d}\xi\;.$
Hence, since $d_{\rho}^{\prime}\leq d_{\rho}$ in $\Omega^{\prime}$,
$\int_{\Omega}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi=\int_{\Omega^{\prime}}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big(\frac{p-1}{p}\Big)^{p}\int_{\Omega^{\prime}}\frac{|u|^{p}}{d_{\rho}^{\prime p}}{\rm d}\xi\geq\Big(\frac{p-1}{p}\Big)^{p}\int_{\Omega}\frac{|u|^{p}}{d_{\rho}^{p}}{\rm d}\xi$
@@
We point that the fact that the key property that $d_{\rho}^{\prime}\leq d_{\rho}$, for $d$ and $d^{\prime}$ as in the proof of Theorem 19 in the bounded convex polytope $\Omega^{\prime}$ and reflects the geometric nature of the Carnot-Carathéodory distance, meaning in particular that that the latter is a “true” distance respecting the geometry of ${\mathbb{H}}^{n}$.
@@ 4. Hardy inequalities on stratified groups of step two
where the $B^{(i)}$’s are $m\times m$ matrices, and $\langle\cdot,\cdot\rangle$ stands for the standard inner product in $\mathbb{R}^{m}$, see e.g. [BLU07, Remark 17.3.1]. The group law (51) can also be written as
@@
where $\langle Bg^{(1)},g^{\prime(1)}\rangle$ denotes the (nm)(n-m)-tuple
@@ 4.1. On the distance function from the boundary of bounded convex domains in stratified groups
Let $\mathbb{H}^{n}$ be the Heisenberg group and $\Pi_{0}=\{(x,y,t)\in\mathbb{H}^{n}:t>0\}$. We shall prove that the the distance $d_{N}$ (cf. (13)) is not weakly $H$-concave.
@@ 4.2. Hardy inequalities with respect to the Euclidean distance
In this section we prove that the Euclidean distance to the boundary on a convex, bounded domain $\Omega$ is weakly $H$-concave and superharmonic. This provides a proof of the $L^{p}$-Hardy inequality for such domains.
@@
From Theorem 23 we obtain the following result; the sharpness of the constant follows under the hypotheses of part (b) of Theorem 1.
@@
Let $G$ be a stratified group of step two and let $\Omega\subset G$ be a bounded domain which is convex in the Euclidean sense. Then for any $p>1$ we have
${\rm(i)}$ $\quad\Delta_{p,H}d\leq 0\;\mbox{ in the distributional sense in }\Omega;$
@@
$\hskip 42.67912pt\int_{\Omega}|\nabla_{H}u|^{p}{\rm d}g\geq\Big(\frac{p-1}{p}\Big)^{p}\int_{\Omega}\frac{|\nabla_{H}d|^{p}}{d^{p}}|u|^{p}{\rm d}g\;,\quad u\in{C^{\infty}_{c}}(\Omega),$
@@
Let $\rho\in C_{c}^{\infty}(G)$ with $\rho\geq 0$ and $\int_{G}\rho\,dx=1$. For $\epsilon>0$ we define
$\rho_{\varepsilon}(x)=\varepsilon^{-Q}\rho(\delta_{\varepsilon^{-1}}x)\,.$
and
$d_{\varepsilon}(x)=(d*\rho_{\varepsilon})(x)=\int_{G}d(xy^{-1})\rho_{\varepsilon}(y)dy\,.$
We will show that $d_{\varepsilon}$ is weakly $H$-concave on the set
$\Omega_{\varepsilon}:=\{x\in\Omega\,:\,B_{CC}(x,\varepsilon)\subset\Omega\}\,.$
Let $g\in\Omega_{\varepsilon}$ and let $g^{\prime}\in H_{g}\cap\Omega_{\epsilon}$. Since $g^{\prime}\in H_{g}$ there exists $v\in V_{1}$ such that
$g^{\prime}=g\exp(v)\in\Omega_{\varepsilon},$
and for every $\lambda\in[0,1]$ we have
$g_{\lambda}:=g\,\delta_{\lambda}(g^{-1}g^{\prime})=g\exp(\lambda v)\,$$\qquad$
implying that the weak $H$-concavity condition can also be written as
$d(g\exp(\lambda v))\geq(1-\lambda)d(g)+\lambda d(g\exp v),\qquad\lambda\in[0,1]\,$,
For $y\in\text{supp}\rho_{\varepsilon}\subset B_{CC}(e,\varepsilon)$, we define the curve
$\gamma_{y}(t):=g\exp(tv)y^{-1},\qquad t\in[0,1]\,.$
Then $\gamma_{y}$ is a horizontal curve inside $\Omega$ connecting $gy^{-1}\in B_{CC}(g,\varepsilon)\subset\Omega$ with $g^{\prime}y^{-1}B_{CC}(g^{\prime},\varepsilon)\subset\Omega$. So
$d(\gamma_{y}(\lambda))\geq(1-\lambda)d(\gamma_{y}(0))+\lambda d(\gamma_{y}(1))\,$,
that is
$d(g\exp(\lambda v)y^{-1})\geq(1-\lambda)d(gy^{-1})+\lambda d(g\exp(v)y^{-1})\,.$
Multiplying by $\rho_{\varepsilon}(y)\geq 0$ and integrating we obtain
$d_{\varepsilon}(g\exp(\lambda v))\geq(1-\lambda)d_{\varepsilon}(g)+\lambda d_{\varepsilon}(g\exp v)\,$,
i.e. $d_{\varepsilon}$ is weakly $H$-concave in $\Omega_{\varepsilon}$.
Now, we have
$\Delta_{p,H}d_{\varepsilon}$ $=\operatorname{div}_{H}\bigl(|\nabla_{H}d_{\varepsilon}|^{p-2}\nabla_{H}d_{\varepsilon}\bigr)$
(56) $=|\nabla_{H}d_{\varepsilon}|^{p-4}\Big(|\nabla_{H}d_{\varepsilon}|^{2}\Delta_{H}d_{\varepsilon}+(p-2)\bigl\langle D^{2}_{H}d_{\varepsilon}\,\nabla_{H}d_{\varepsilon},\nabla_{H}d_{\varepsilon}\bigr\rangle\Big),$
where
$D^{2}_{H}d_{\varepsilon}:=\left(\frac{1}{2}\bigl(X_{i}X_{j}d_{\varepsilon}+X_{j}X_{i}d_{\varepsilon}\bigr)\right)_{i,j=1}^{2n}$
is the symmetrized horizontal Hessian of $d_{\epsilon}$ (here $X_{n+k}=Y_{k}$). By [DGN03, Theorem 5.12] $D_{H}^{2}d_{\varepsilon}$ is negative semidefinite on $\Omega_{\varepsilon}$. Using also the elementary matrix inequality $A\leq{\rm tr}(A)I_{d}$ which is valid from any symmetric positive definite matrix, we conclude from (56) that
$\Delta_{p,H}d_{\varepsilon}\leq 0\,,\qquad\text{pointwise on }\Omega_{\varepsilon}\,.$
Finally, to show that $\Delta_{p,H}d\leq 0$ in the weak sense we need to show that for any $\varphi\in C_{c}^{\infty}(\Omega)$
(57) $\lim_{\varepsilon\to 0}\int_{\Omega}|\nabla_{H}d_{\varepsilon}|^{p-2}\langle\nabla_{H}d_{\varepsilon},\nabla_{H}\varphi\rangle dx=\int_{\Omega}|\nabla_{H}d|^{p-2}\langle\nabla_{H}d,\nabla_{H}\varphi\rangle dx\,.$
Since
$\nabla_{H}d_{\varepsilon}\to\nabla_{H}d\,,\qquad\text{a.e. in supp}(\varphi)\,$,
we have
$\langle|\nabla_{H}d_{\varepsilon}|^{p-2}\nabla_{H}d_{\varepsilon},\nabla_{H}\varphi\rangle\to\langle|\nabla_{H}d|^{p-2}\nabla_{H}d\,\nabla_{H}\varphi\rangle$
Now, notice that since $|\nabla_{h}d_{\varepsilon}|$ is bounded on ${\rm supp}(\varphi)$ we have
$\left|\langle|\nabla_{H}d_{\varepsilon}|^{p-2}\nabla_{H}d_{\varepsilon},\nabla_{H}\varphi\rangle\right|\leq C_{\varphi}\,$,
and we can apply the dominated convergence theorem to get
$\lim_{\varepsilon\to 0}\langle\Delta_{p,H}d_{\varepsilon},\varphi\rangle=-\int_{\Omega}|\nabla_{H}d|^{p-2}\langle\nabla_{H}d,\nabla_{H}\varphi\rangle\,dg=\langle\Delta_{p,H}d,\varphi\rangle\,$,
which completes the proof of (i) Part (ii) then follows from Theorem 1.
Appendix
We prove here the following result used in the proof of Theorem 11.
Proposition 25.
Let $\Pi_{1}$, $\Pi_{2}$ be two half-spaces in the Heisenberg group $\mathbb{H}^{n}$ and let $d_{N,1}$, $d_{N,2}$ denote the correspoding distances to the boundary with respect to the gauge quasi-norm $N$, cf. (13). If there exists $\xi\in\mathbb{H}^{n}$ such that
[…]
then $\Pi_{1}=\Pi_{2}$.
Proof.
By group action there exist $\xi_{1},\xi_{2}\in\mathbb{H}^{n}$ such that
$d_{N,k}(\xi)={\rm dist}_{N}(\xi_{k}\,\xi,\partial\Pi_{0})=:d_{N}(\xi_{k}\,\xi),\quad k=1,2,\;\;\xi\in\Pi_{k},$
where $\Pi_{0}$ and $d_{N}$ is as in Lemma 6.
Given $\xi\in\Pi_{1}\cap\Pi_{2}$ we set
$\xi_{k}\xi=:\eta_{k}=:(\alpha_{k},\beta_{k},\tau_{k})\;,\qquad k=1,2.$
By Lemma 6 $d_{N}(\eta_{k})$ depends only on $\rho_{k}:=\sqrt{|\alpha_{k}|^{2}+|\beta_{k}|^{2}}$ and $\tau_{k}$. Let $s_{k}$, $k=1,2$, be defined by
(58) $s_{k}^{3}+2s_{k}=\frac{\tau_{k}}{\rho_{k}^{2}}.$
If $d_{N,1}(\xi)=d_{N,2}(\xi)$ then $d_{N}(\eta_{1})=d_{N}(\eta_{2})$ and recalling also Remark 3 we obtain
(59) $\rho_{1}^{4}\,s_{1}^{4}\,(s_{1}^{2}+1)=\rho_{2}^{4}\,s_{2}^{4}\,(s_{2}^{2}+1).$
Writing $\xi=(x,y,t)$ and $\xi_{k}=(x_{k},y_{k},t_{k})$ we have
$\eta_{k}=\big(x_{k}+x,y_{k}+y,t_{k}+t+2(x_{k}\cdot y-x\cdot y_{k})\big).$
Assume now that we additionally have $\nabla d_{N,1}(\xi)=\nabla d_{N,2}(\xi)$. There holds
$\frac{\partial{d_{N,k}}}{\partial{t}}(\xi)=\frac{\partial{d_{N}}}{\partial{\tau_{k}}}(\eta)\,.$
Hence, using again Remark 3,
(60) $\rho_{1}^{2}\,s_{1}^{3}=\rho_{2}^{2}\,s_{2}^{3}.$
Combining (59) and (60) and recalling that $s_{1}$ and $s_{2}$ are positive we obtain that
(61) $s_{1}=s_{2}=:s\;,\qquad\qquad\rho_{1}=\rho_{2}=:\rho.$
Now, since $d_{N}(\eta_{k})$ depends only on $\rho_{k}$ and $\tau_{k}$, we have by Remark 3 that
$\frac{\partial{d_{N,k}}}{\partial{x_{i}}}(\xi)$ $=\frac{\partial{d_{N}}}{\partial{\rho_{k}}}\frac{\partial{\rho_{k}}}{\partial{x_{i}}}+\frac{\partial{d_{N}}}{\partial{\tau_{k}}}\frac{\partial{tau_{k}}}{\partial{x_{i}}}$
$=\frac{1}{\rho_{k}}\frac{\partial{d_{N}}}{\partial{\rho_{k}}}(x_{i}+x_{k,i})-2\frac{\partial{d_{N}}}{\partial{\tau_{k}}}y_{k,i}$
$=-\frac{\rho_{k}^{2}s_{k}^{4}}{d_{N,k}^{3}}(x_{i}+x_{k,i})-\frac{\rho_{k}^{2}s_{k}^{3}}{2d_{N,k}^{3}}y_{k,i},$
and similarly
$\frac{\partial{d_{N,k}}}{\partial{y_{i}}}(\xi)=-\frac{\rho_{k}^{2}s_{k}^{4}}{d_{N,k}^{3}}(y_{i}+y_{k,i})+\frac{\rho_{k}^{2}s_{k}^{3}}{2d_{N,k}^{3}}x_{k,i}.$
Recalling also (61), it follows that at a point $\xi$ as in the statement we have
$-s(x_{i}+x_{1,i})-\frac{y_{1,i}}{2}=-s(x_{i}+x_{2,i})-\frac{y_{2,i}}{2}$
$-s(y_{i}+y_{1,i})+\frac{x_{1,i}}{2}=-s(y_{i}+y_{2,i})+\frac{x_{2,i}}{2}$
and therefore $x_{1}=x_{2}$ and $y_{1}=y_{2}$. Finally, from (58) we also obtain $\tau_{1}=\tau_{2}$ and therefore $t_{1}=t_{2}$. Hence $\xi_{1}=\xi_{2}$ as required. ∎
Acknowledgment. We would like to thank the two anonymous referees for the careful reading of the paper and for their useful comments.